Daniel and Jan agreed to pay $566,000 for a four-bedroom colonial home in Waltham, Massachusetts, with a $60,000 down payment. They have a 30-year mortgage at a fixed rate of 6.00%.

  
a. How much is their monthly payment? (Do not round intermediate calculations. Round your answer to the nearest cent.)

b. After the first payment, what would be the balance of the principal?(Do not round intermediate calculations. Round your answers to the nearest cent.)
Answer #1

a) Monthly payment:-

=PMT(rate,nper,pv)

=PMT(6%/12,30*12,566000-60000)

=3033.73

b) Balance principal after first payment:-

Interest paid =(566000-60000)*6%/12 =2530

Principal paid =3033.73-2530 =503.73